OpenStudy (anonymous):

Let F(x) = - 2x+ 2, find F (1over2) =

jimthompson5910 (jim_thompson5910):

replace x with 1/2 and evaluate

OpenStudy (anonymous):

Just change x to \[\frac{ 1 }{ 2 }\]. You will have: \[-2*(\frac{ 1 }{ 2 })+2 = -1 +2 = 1\]

OpenStudy (anonymous):

ok thanks